# Discovering mathematical concepts through a multi-agent system

### 04 Mar 2026 v1

@@ 1.7 Learning problem and mathematical background
Consider the vector space over a field $k$ (with characteristic $\neq 2$) freely generated by the $i$-simplices (simplices of dimension $i$); denote this space by $C_{i}$. We define the maps $\partial_{i}:C_{i}\rightarrow C_{i-1}$ by sending a simplex to the sum of simplices on its boundary, where the coefficients $\pm 1$ respect orientations. It is easy to show that these maps satisfy $\partial_{i-1}\circ\partial_{i}=0$, an important property for the definitions that follow131313‘The boundary of a boundary is zero’ [Wheeler1989-WHEIPQ]..
@@
For a connected surface, $b_{2}$ is $1$ if and only if the surface is orientable.
@@ 1.8 Reinforcement learning
The fundamental loop involves an agent choosing an action $a$ in a state $s$, which transitions it to a new state $s^{\prime}$ via the probability $p(s^{\prime}|s,a)$, resulting in a reward $r(s,a,s^{\prime})$ [^17] This can also be probabilistic.. All of this is strictly determined by $M$. For example, in chess $S$ is naturally taken to be the space of board positions and $A$ individual moves. Then $P$ is deterministic and $R$ is only allocated at the end of games. Given an MDP, the sequence $\{s_{0},s_{1},..\}$ is a Markov process. We call the position within this the timestep $t$.
@@
MDPs find applications in economics and other areas. Under certain assumptions [^18] For example, finite state and action spaces and bounded rewards., one can compute optimal policies with linear programming, assuming full knowledge of the MDP. However, in complex settings, where transitions are unknown and the $S$ and $A$ may be infinite or continuous, data-driven approximation techniques become necessary. This is really what is referred to as reinforcement learning. For instance, one can parameterise policies and expectations of future reward using neural networks.
@@ 1.10 Conjecturing agent
we first consider the data-dependent term $\mathcal{L}_{\mathcal{D}}$. The expression is true or false on a datapoint $d_{i}$, so $\tilde{s}(d_{i})$ is a Boolean quantity. For example, Eq. 3 will evaluate to $1$ on topological spheres and 0 on topological tori. The weighted accuracy and full data-dependent terms are:
@@ 1.12 Environment and provability
The provability score models the proving side of mathematics, assigning $\rho(s)$ to measure how close $s$ is to being ‘provable’ with the system’s knowledge. In general this may not be well defined, but we can consider simple approximations. Given an automated prover $A$, let:
@@
One choice for $A$ is an LLM prover, e.g. Lean Copilot [song_lean_2025]. This attempts to prove the posed statements using a set of relevant theorems as background, such as the rank-nullity theorems for the $C_{i}$. For translation into Lean, we used a handcrafted code, with a header and premise set containing what we wanted $A$ to ‘know’; this prevented any hallucination effects caused by autoformalisation [wu2022autoformalization].
Since our experiments mostly took place within linear algebra, we were also able to prove any correct statements with a prewritten proof, in which case the system is LLM free [^27] This was done through a combination of tactics including $\mathtt{grind}$ and $\mathtt{ring}$.. The tests produced the same qualitative results with either approach. We also tried some simple alternatives to $A$, discussed further in §2.2.
For a statement to obtain any reward or trigger termination, it must pass simple non-degeneracy checks. As an example, we discourage formulas that are true on all of $\mathcal{D}$, adding those that are noticed to the premises [^28] Such as the rank-nullity theorems.. These are somewhat discretionary choices, but the relationship between the knowledge of $A$ and the data should have a large impact on performance.
@@ 2.1 Premises and data
For $\mathcal{D}_{0}$ we include $b_{0}=1$ and $b_{2}=1$ (representing connectedness and orientability) in $\mathcal{P}$, and just search for statements like $\{{b_{1}=0}\implies\chi=2\}$. For $\mathcal{D}_{2}$ the only premises are rank-nullity theorems, so we may hope for a conjecture similar to Eq. 6. Even if premises involving $b_{i}$ are used by $A$, the only effect is through $\rho$ feedback. In a new problem, finding an appropriate relationship between $\mathcal{P}$ and $\mathcal{D}$ may require experimentation.
@@ 2.2 Different models
  • $\mathcal{M}_{0}$ - Full System: The system with all agents and environmental components. In $\mathcal{M}_{0}:\operatorname{\bm{\mathtt{CA}}}+\operatorname{\bm{\mathtt{SA}}}+$Provability, we use the usual automated prover $A$ that assigns the discrete score $\rho(s)$ (Eq. 5). In $\mathcal{M}_{0}+$Noise : $\operatorname{\bm{\mathtt{CA}}}+\operatorname{\bm{\mathtt{SA}}}+$Noisy Provability, we use a stochastic variant, adding Gaussian noise to $\rho(s)$ and reducing the provability threshold. This is a basic test for how the system performs in a more uncertain, general setting.
@@ 3 Discussion
Acknowledgements The authors thank Minhyong Kim for many helpful discussions. Daattavya Aggarwal also thanks Anand Rao Tadipatri for suggestions on using Lean. Oisin Kim was supported by the Accelerate Programme for Scientific Discovery.

### 29 Mar 2026 v2

@@ 1.7 Learning problem and mathematical background
Consider the vector space over a field $k$ (with characteristic $\neq 2$) freely generated by the $i$-simplices (simplices of dimension ii); denote this space by $C_{i}$. We define the maps $\partial_{i}:C_{i}\rightarrow C_{i-1}$ by sending a simplex to the sum of simplices on its boundary, where the coefficients $\pm 1$ respect orientations. It is easy to show that these maps satisfy $\partial_{i-1}\circ\partial_{i}=0$, an important property for the definitions that follow131313‘The boundary of a boundary is zero’ [Wheeler1989-WHEIPQ]..
@@
For a connected surface, $b_{2}$ is 11 if and only if the surface is orientable.
@@ 1.8 Reinforcement learning
The fundamental loop involves an agent choosing an action $a$ in a state $s$, which transitions it to a new state $s^{\prime}$ via the probability $p(s^{\prime}|s,a)$, resulting in a reward $r(s,a,s^{\prime})$ [^17] This can also be probabilistic.. All of this is strictly determined by $M$. For example, in chess $S$ is naturally taken to be the space of board positions and AA individual moves. Then $P$ is deterministic and $R$ is only allocated at the end of games. Given an MDP, the sequence $\{s_{0},s_{1},..\}$ is a Markov process. We call the position within this the timestep $t$.
@@
MDPs find applications in economics and other areas. Under certain assumptions [^18] For example, finite state and action spaces and bounded rewards., one can compute optimal policies with linear programming, assuming full knowledge of the MDP. However, in complex settings, where transitions are unknown and the $S$ and AA may be infinite or continuous, data-driven approximation techniques become necessary. This is really what is referred to as reinforcement learning. For instance, one can parameterise policies and expectations of future reward using neural networks.
@@ 1.10 Conjecturing agent
we first consider the data-dependent term $\mathcal{L}_{\mathcal{D}}$. The expression is true or false on a datapoint $d_{i}$, so $\tilde{s}(d_{i})$ is a Boolean quantity. For example, Eq. 3 will evaluate to 11 on topological spheres and 0 on topological tori. The weighted accuracy and full data-dependent terms are:
@@ 1.12 Environment and provability
The provability score models the proving side of mathematics, assigning $\rho(s)$ to measure how close $s$ is to being ‘provable’ with the system’s knowledge. In general this may not be well defined, but we can consider simple approximations. Given an automated prover AA, let:
@@
One choice for AA is an LLM prover, e.g. Lean Copilot [song_lean_2025]. This attempts to prove the posed statements using a set of relevant theorems as background, such as the rank-nullity theorems for the $C_{i}$. For translation into Lean, we used a handcrafted code, with a header and premise set containing what we wanted AA to ‘know’; this prevented any hallucination effects caused by autoformalisation [wu2022autoformalization].
Since our experiments mostly took place within linear algebra, we were also able to prove any correct statements with a prewritten proof, in which case the system is LLM free [^27] This was done through a combination of tactics including $\mathtt{grind}$ and $\mathtt{ring}$.. The tests produced the same qualitative results with either approach. We also tried some simple alternatives to AA, discussed further in §$2.$
For a statement to obtain any reward or trigger termination, it must pass simple non-degeneracy checks. As an example, we discourage formulas that are true on all of $\mathcal{D}$, adding those that are noticed to the premises [^28] Such as the rank-nullity theorems.. These are somewhat discretionary choices, but the relationship between the knowledge of AA and the data should have a large impact on performance.
@@ 2.1 Premises and data
For $\mathcal{D}_{0}$ we include $b_{0}=1$ and $b_{2}=1$ (representing connectedness and orientability) in $\mathcal{P}$, and just search for statements like $\{{b_{1}=0}\implies\chi=2\}$. For $\mathcal{D}_{2}$ the only premises are rank-nullity theorems, so we may hope for a conjecture similar to Eq. 6. Even if premises involving $b_{i}$ are used by AA, the only effect is through $\rho$ feedback. In a new problem, finding an appropriate relationship between $\mathcal{P}$ and $\mathcal{D}$ may require experimentation.
@@ 2.2 Different models
  • $\mathcal{M}_{0}$ - Full System: The system with all agents and environmental components. In $\mathcal{M}_{0}:\operatorname{\bm{\mathtt{CA}}}+\operatorname{\bm{\mathtt{SA}}}+$Provability, we use the usual automated prover AA that assigns the discrete score $\rho(s)$ (Eq. 5). In $\mathcal{M}_{0}+$Noise : $\operatorname{\bm{\mathtt{CA}}}+\operatorname{\bm{\mathtt{SA}}}+$Noisy Provability, we use a stochastic variant, adding Gaussian noise to $\rho(s)$ and reducing the provability threshold. This is a basic test for how the system performs in a more uncertain, general setting.
@@ 3 Discussion
Acknowledgements: The authors thank Minhyong Kim for many helpful discussions. Daattavya Aggarwal also thanks Anand Rao Tadipatri for suggestions on using Lean. Oisin Kim was supported by the Accelerate Programme for Scientific Discovery.
Code Availability: The full code along with trained models is available on the repository: https://github.com/daattavya98/MathWorld-learning-homology.